# Convergence to equilibrium for a degenerate three species reaction-diffusion system

### v1

@@ 1. Introduction
(12) {$D(a,b,c)\geq&\frac{d_{a}}{C(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{d_{c}}{C(\Omega)}\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{L^{2}(\Omega)}\qquad\mbox{ for }\,d_{b}=0,\\[5.69046pt] D(a,b,c)\geq&\frac{d_{a}}{C(\Omega)}\left\|\delta_{A}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{d_{b}}{C(\Omega)}\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{L^{2}(\Omega)}\qquad\mbox{ for }\,d_{c}=0.\end{aligned}\right.$
An application of Poincaré-Writinger inequality (see Lemma 7 in the Appendix) for the dimension  $N\geq 4$ gives us for $d_{b}=0$,
$D(a,b,c)\geq\frac{d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\frac{d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
and the same inequality for dimension $N=1,2,3$, gives us for $d_{b}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{6}(\Omega)}+\frac{4d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{L^{6}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
where $P(\Omega)$ is the Poincaré-Writinger constant (see Lemma 7 in the Appendix).
A similar application of Poincaré-Writinger inequality for $N\geq 4$ gives us for $d_{c}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{L^{\frac{2N}{N-2}}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
and the same inequality for the dimension $N=1,2,3$, gives us for $d_{c}=0$,
$D(a,b,c)\geq\frac{4d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{L^{6}(\Omega)}+\frac{4d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{L^{6}(\Omega)}+\|AB-C\|^{2}_{L^{2}(\Omega)}$
where $P(\Omega)$ is the Poincaré-Writinger constant (see Lemma 7 in the Appendix)
Furthermore for both the cases we have the following relation between the two functional by the time derivative i.e
$\frac{d}{dt}(E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty}))=-D(a,b,c)$
Furthermore consider the function $\Gamma(x,y)=\frac{x\ln{\frac{x}{y}}-x+y}{(\sqrt{x}-\sqrt{y})^{2}}$ ; we see that we can find a Constant $C_{\Gamma}>0$; such that $\Gamma(x,y)\leq C_{\Gamma}\max\{1,\ln{\frac{x}{y}}\}$. We are going to represent difference between entropy by this function. We can write:
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$
$=\int_{\Omega}\Gamma(a,a_{\infty})(A-A_{\infty})^{2}$ $+\int_{\Omega}\Gamma(b,b_{\infty})(B-B_{\infty})^{2}+\int_{\Omega}\Gamma(c,c_{\infty})(C-C_{\infty})^{2}$
$\leq C_{\Gamma}\max\big{\{}1,\ln{(\|a\|_{L^{\infty}(\Omega)}+1)}+$ $|\ln{a_{\infty}}|\big{\}}\|A-A_{\infty}\|_{L^{2}(\Omega)}^{2}$
$+C_{\Gamma}\max\big{\{}1,\ln{(\|b\|_{L^{\infty}(\Omega)}+1)}$ $+|\ln{b_{\infty}}|\big{\}}\|B-B_{\infty}\|_{L^{2}(\Omega)}^{2}$
$+C_{\Gamma}\max\big{\{}1,\ln{(\|c\|_{L^{\infty}(\Omega)}+1)}$ $+|\ln{c_{\infty}}|\big{\}}\|C-C_{\infty}\|_{L^{2}(\Omega)}^{2}$
Now for degeneracy $d_{b}=0$, we derive in Lemma 4 and for degeneracy $d_{c}=0$, we derive in Proposition 8 that the growth of the species varies polynomially with time, which leads us to the following relation: for some positive constant $C_{1}>0$, and time sufficiently large time i.e $t\geq T_{\epsilon}$, where $\epsilon>0$,
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$
$\leq$ $C_{1}(1+t)^{\epsilon}\Big{(}\|A-A_{\infty}\|_{L^{2}(\Omega)}^{2}+\|B-B_{\infty}\|_{L^{2}(\Omega)}^{2}+\|C-C_{\infty}\|_{L^{2}(\Omega)}^{2}\Big{)}$
@@
Now we have from [DF06]; $\exists C_{EB}>0$, depending only on the domain and equilibrium point such that:
$\|\sqrt{\overline{A^{2}}}-A_{\infty}$ $\|_{L^{2}(\Omega)}^{2}+\|\sqrt{\overline{B^{2}}}-B_{\infty}\|_{L^{2}(\Omega)}^{2}+\|\sqrt{\overline{C^{2}}}-C_{\infty}\|_{L^{2}(\Omega)}^{2}$
$\leq$ $C_{EB}\big{(}\|\delta_{A}\|_{L^{2}(\Omega)}^{2}+\|\delta_{B}\|_{L^{2}(\Omega)}^{2}+\|\delta_{C}\|_{L^{2}(\Omega)}^{2}+\|C-AB\|_{L^{2}(\Omega)}^{2}\big{)}$
@@
This indicates that we will able to apply Grónwall inequality if we somehow link entropy dissipation with the missing term $\|\delta_{B}\|^{2}_{L^{2}(\Omega)}$ for $d_{b}=0$ case and similarly with the missing term $\|\delta_{C}\|^{2}_{L^{2}(\Omega)}$ for $d_{c}=0$ case. So to obtain exponential decay in convergence type of results our goal is to relate entropy dissipation with the missing term.
We will relate the missing term with entropy dissipation in Proposition 4 for $d_{b}=0$ case in dimension $N\geq 4$, and proved the same for $d_{b}=0$ case in dimension $N=1,2,3$ in Proposition 5 in the following way:
@@
To prove this we require an estimate of $L^{\frac{N}{2}}(\Omega)$ norm of $b$ for $N\geq 4$ and an estimate of $L^{\frac{3}{2}}(\Omega)$ norm of $b$ for $N=1,2,3$. These estimates are the Key Estimates of our article. It turns out our Key estimates are the following:
  • For $N\geq 4$   $\|b\|_{L^{\frac{N}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{N-2}{N-1}}$
  • For $N=1,2,3;\ \ \|b\|_{L^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}$
[…]
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq C_{0}\qquad\forall\tau\geq 0$
This above result is proved in Lemma 1, and this will help us to prove our key estimate for dimension $N\geq 4$. Similarly we will able to prove time dependent evolution of space integral of diffusive species for degenerate $d_{b}=0$ case in dimension $N=1,2,3$, in Proposition 3 and Lemma 3. The relation for dimension $N=1,2,3$ for $d_{b}=0$ case is described below:for some natural number $\mu_{C}>0$,
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{c}(1+\tau)^{\mu_{C}}\qquad\forall\tau\geq 0$
In this article our calculations are inspired from the article [EMT20].
@@
For $d_{c}=0$, we will relate entropy dissipation with $\delta_{C}$. In this article for $d_{c}=0$ we consider the dimension $N\leq 4$. Although we will establish our result only for dimension $N=3$, for dimension $N=1,2,4$ all the calculations are similar. In Proposition 9 We establish the following relations between entropy dissipation and $\delta_{C}$:
$D(a,b,c)\geq\hat{C}(1+t)^{-\frac{1}{3}}\big{(}\|A-\overline{A}\|_{L^{2}(\Omega)}^{2}+\|B-\overline{B}\|_{L^{2}(\Omega)}^{2}+\|C-\overline{C}\|_{L^{2}(\Omega)}^{2}\big{)}$
@@
To prove this we need a particular integral estimate of species $a$, which we will establish in Lemma 6. The particular relation is the following:
$\|a\|_{L^{\frac{3}{2}}(\Omega)}\leq\tilde{C}(1+t)^{\frac{1}{3}};\qquad\forall t\geq 0.$
@@
This is the key estimate for the degenerate case $d_{c}=0$; to establish this we will use the following estimate (see Lemma 5): $\exists C_{\Omega}>0$ independent of time such that;
$\int_{0}^{T}\int_{\Omega}a^{2}\leq C_{\Omega}(1+T);\ \ \int_{0}^{T}\int_{\Omega}b^{2}\leq C_{\Omega}(1+T)\qquad\forall T\geq 0$
@@
These results along with Grónwall inequality will help us to show entropy functional for both degeneracy cases $d_{b}=0$ and $d_{c}=0$ decays sub-exponentially with time. On the other hand an application of Cziszár-Kullback-Pinsker Inequality provides us the following lower bound of entropy functional; [DF06]
$E(a,b,c)$ $-E(a_{\infty},b_{\infty},c_{\infty})\geq\frac{(3+2\sqrt{2})|\Omega|}{2M_{1}(9+2\sqrt{2})}\|a-a_{\infty}\|_{L^{1}(\Omega)}^{2}$
$+$ $\frac{(3+2\sqrt{2})|\Omega|}{2M_{2}(9+2\sqrt{2})}\|b-b_{\infty}\|_{L^{1}(\Omega)}^{2}+\frac{(3+2\sqrt{2})|\Omega|}{(M_{1}+M_{2})(9+2\sqrt{2})}\|c-c_{\infty}\|_{L^{1}(\Omega)}^{2}$
Combining all these we arrive at the main results of our article i.e Theorem 1 and Theorem 2.

### v2

@@ 1. Introduction
(12) {$D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{4\,d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)},\qquad\mbox{ for }\,d_{b}=0,\\ D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{4\,d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)},\qquad\mbox{ for }\,d_{c}=0,\\ \end{aligned}\right.$
where $P(\Omega)$ is the Poincaré constant of the domain $\Omega$ (see Lemma 7 in the Appendix). Observe that the deviation term $\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing in the above lower bound corresponding to the degenerate case $d_{b}=0$ and the deviation term $\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing in the lower bound corresponding to the degenerate case $d_{c}=0$. Note that, for the dissipation functional of the degenerate system (2), the Poincaré inequality actually yields
{$D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\frac{4\,d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)}\qquad\mbox{ for }\,N\geq 4,\\ D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{4\,d_{c}}{P(\Omega)}\|\delta_{C}\|^{2}_{\mathrm{L}^{6}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)}\qquad\qquad\quad\mbox{ for }N\leq 3.\end{aligned}\right.$
Similarly, for the dissipation functional of the degenerate system (3), we have
{$D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+\frac{4\,d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{\mathrm{L}^{\frac{2N}{N-2}}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)}\qquad\mbox{ for }N\geq 4,\\ D(a,b,c)\geq&\frac{4\,d_{a}}{P(\Omega)}\|\delta_{A}\|^{2}_{\mathrm{L}^{6}(\Omega)}+\frac{4\,d_{b}}{P(\Omega)}\|\delta_{B}\|^{2}_{\mathrm{L}^{6}(\Omega)}+4\|AB-C\|^{2}_{\mathrm{L}^{2}(\Omega)}\qquad\qquad\quad\mbox{ for }N\leq 3.\end{aligned}\right.$
Note that, if $(a,b,c)$ solves either of the system (2) or (3), we have
$\frac{d}{dt}\Big{(}E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\Big{)}=-D(a,b,c),$
i.e. the relative entropy (relative with respect to the homogeneous equilibrium state) is non-increasing in time. The relative entropy has the following expression:
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})=\int_{\Omega}\big{(}a\ln{a}-a-a_{\infty}$ $\ln{a_{\infty}}+a_{\infty}\big{)}\,{\rm d}x+\int_{\Omega}\big{(}b\ln{b}-b-b_{\infty}\ln{b_{\infty}}+b_{\infty}\big{)}\,{\rm d}x$
$+\int_{\Omega}\big{(}c\ln{c}-c-c_{\infty}\ln{c_{\infty}}+c_{\infty}\big{)}\,{\rm d}x.$
The above expression rewrites as
(13) $E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})=\int_{\Omega}\left(a\ln{\left(\frac{a}{a_{\infty}}\right)}-a+a_{\infty}\right)\,{\rm d}x+\int_{\Omega}$ $\left(b\ln{\left(\frac{b}{b_{\infty}}\right)}-b+b_{\infty}\right)\,{\rm d}x$
$+\int_{\Omega}\left(c\ln{\left(\frac{c}{c_{\infty}}\right)}-c+c_{\infty}\right)\,{\rm d}x.$
Define a function $\Gamma:(0,\infty)\times(0,\infty)\to\mathbb{R}$ as follows:
(14) $\Gamma(x,y):=\left\{\begin{aligned} &\frac{x\ln\left(\frac{x}{y}\right)-x+y}{\left(\sqrt{x}-\sqrt{y}\right)^{2}}\qquad\mbox{ for }x\not=y,\\ &2\qquad\qquad\qquad\qquad\ \ \ \mbox{ for }x=y.\end{aligned}\right.$
It can be shown (see [4, Lemma 2.1, p.162] for details) that the above defined function satisfies the following bound:
$\Gamma(x,y)\leq C_{\Gamma}\max\left\{1,\ln\left(\frac{x}{y}\right)\right\}$
for some positive constant $C_{\Gamma}$. Note that using the function $\Gamma$ defined above, the relative entropy can be rewritten as
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})=\int_{\Omega}\Gamma(a,a_{\infty})(A-A_{\infty})^{2}+\int_{\Omega}\Gamma(b,b_{\infty})(B-B_{\infty})^{2}+\int_{\Omega}\Gamma(c,c_{\infty})(C-C_{\infty})^{2}.$
Using the aforementioned bound for $\Gamma$, we obtain
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq C_{\Gamma}$ $\max\big{\{}1,\ln{(\|a\|_{\mathrm{L}^{\infty}(\Omega)}+1)}+|\ln{a_{\infty}}|\big{\}}\|A-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}$
$+$ $C_{\Gamma}\max\big{\{}1,\ln{(\|b\|_{\mathrm{L}^{\infty}(\Omega)}+1)}+|\ln{b_{\infty}}|\big{\}}\|B-B_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}$
$+$ $C_{\Gamma}\max\big{\{}1,\ln{(\|c\|_{\mathrm{L}^{\infty}(\Omega)}+1)}+|\ln{c_{\infty}}|\big{\}}\|C-C_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}.$
We derive the growth rate of the solutions corresponding to (2) and (3) in Lemma 4 and in Proposition 8 respectively. Both these results assert that the growth of the solutions can at most be polynomial in time. This leads to
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})\leq C_{1}(1+t)^{\varepsilon}\Big{(}\|A-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|B-B_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|C-C_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\Big{)},$
for all $t\geq T_{\varepsilon}$, where $\varepsilon\ll 1$ is some positive quantity and $T_{\varepsilon}$ depends on $\varepsilon$. Furthermore, the positive constant $C_{1}$ depends only on the initial data, the domain $\Omega$ and the dimension $N$.
Observe that, other than the logarithm of the growth of the supremum norm of the solution, the growth of relative entropy depends on the $\mathrm{L}^{2}(\Omega)$ norm of the deviation of the $(A,B,C)$ from $(A_{\infty},B_{\infty},C_{\infty})$. Recall from (12) that the dissipation functional is also related to the $\mathrm{L}^{2}(\Omega)$ norm of the deviations $\delta_{A}$, $\delta_{B}$ and $\delta_{C}$. The following observation holds for all the three species. However, we choose to show it for the term $A$:
$\|A-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\leq 3\left(\|A-\overline{A}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|\overline{A}-\sqrt{\overline{A^{2}}}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|\sqrt{\overline{A^{2}}}-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\right).$
The following observation says that the first term on the right hand side dominates the second term:
$\|\overline{A}-\sqrt{\overline{A^{2}}}\|_{\mathrm{L}^{2}(\Omega)}^{2}=$ $|\Omega|\big{|}\overline{A}-\sqrt{\overline{A^{2}}}\big{|}^{2}=|\Omega|\left(\overline{A^{2}}+\overline{A}^{2}-2\overline{A}\sqrt{\overline{A^{2}}}\right)$
$\leq|\Omega|\left(\overline{A^{2}}-\overline{A}^{2}\right)\leq\|A-\overline{A}\|_{\mathrm{L}^{2}(\Omega)}^{2}.$
Here we used the fact that ${\overline{A}\leq\sqrt{\overline{A^{2}}}}$, thanks to Hölder inequality. Hence
$\|A-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\leq 6\left(\|A-\overline{A}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|\sqrt{\overline{A^{2}}}-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\right).$
This helps us deduce
$\|A-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|B-B_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}$ $+\|C-C_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\leq 6\left(\|A-\overline{A}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|B-\overline{B}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|C-\overline{C}\|_{\mathrm{L}^{2}(\Omega)}^{2}\right)$
$+$ $6\left(\|\sqrt{\overline{A^{2}}}-A_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|\sqrt{\overline{B^{2}}}-B_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|\sqrt{\overline{C^{2}}}-C_{\infty}\|_{\mathrm{L}^{2}(\Omega)}^{2}\right).$
@@
Next we borrow a result from [4], which says that there exists $C_{EB}>0$, depending only on the domain and the equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$ such that
(15) $\|\sqrt{\overline{A^{2}}}-A_{\infty}\|_{L^{2}(\Omega)}^{2}+$ $\|\sqrt{\overline{B^{2}}}-B_{\infty}\|_{L^{2}(\Omega)}^{2}+\|\sqrt{\overline{C^{2}}}-C_{\infty}\|_{L^{2}(\Omega)}^{2}$
$\leq$ $C_{EB}\left(\|\delta_{A}\|_{L^{2}(\Omega)}^{2}+\|\delta_{B}\|_{L^{2}(\Omega)}^{2}+\|\delta_{C}\|_{L^{2}(\Omega)}^{2}+\|C-AB\|_{L^{2}(\Omega)}^{2}\right).$
@@
This indicates that to relate entropy with entropy dissipation, we need to relate dissipation with the missing term $\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}$ for the degenerate case (2) and similarly with the missing term $\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}$ for the degenerate case (3).
For the degenerate system (2), the relation between the missing term and the entropy dissipation is shown in Proposition 4 for dimensions $N\geq 4$. For dimensions $N\leq 3$, we obtain the relation in Proposition 5. More precisely, we have obtained
{$D(a,b,c)\geq&\,\hat{C}(1+t)^{-\frac{N-2}{N-1}}\|\delta_{B}\|_{\mathrm{L}^{2}(\Omega)}^{2}\qquad\mbox{ for }N\geq 4,\\ D(a,b,c)\geq&\,\hat{C}(1+t)^{-\frac{5}{6}}\|\delta_{B}\|_{\mathrm{L}^{2}(\Omega)}^{2}\qquad\quad\mbox{ for }N\leq 3.\end{aligned}\right.$
where $\hat{C}$ is some positive constant, independent of time.
In order to obtain the above lower bound, we require an estimate on the growth (in time) of the $\mathrm{L}^{\frac{N}{2}}(\Omega)$ norm of the degenerate species $b$ for dimension $N\geq 4$ and for dimension $N\leq 3$, we require a similar growth estimate of the $\mathrm{L}^{\frac{3}{2}}(\Omega)$ norm of the degenerate species $b$. These estimates are obtained in Lemma 2, Proposition 3 and Lemma 3. More precisely, we have obtained
@@
where $\hat{K}$ is a positive constant independent of time. For dimension $N\leq 3$, the above integral estimation of the species $b$ follows from Gagliardo-Nirenberg inequality whereas for dimension $N\geq 4$, we need closeness assumption (10) on the other two non-zero diffusion coefficients $d_{a}$ and $d_{c}$. Closeness assumption further helps us to estimate $\mathrm{L}^{p}$ integral growth of $a$ and $c$ on a parabolic cylinder with unit height, for some large exponent $p$. It turns out if $d_{a}$ and $d_{c}$ satisfy the closeness condition (10) for dimension $N\geq 4$, then there exists a positive constant $C_{0}>0$, depending on $p$ and independent of time, such that
$\|a\|_{\mathrm{L}^{p}((\tau,\tau+1)\times\Omega)}+\|c\|_{\mathrm{L}^{p}((\tau,\tau+1)\times\Omega)}\leq C_{0}\qquad\forall\tau\geq 0,\ p>N.$
The above result is proved in Lemma 1. For dimension $N\geq 4$, this will lead us to prove our key $\mathrm{L}^{\frac{N}{2}}(\Omega)$ estimate. These strategies are inspired from the article [7], where the authors study existence and large time behaviour of a particular $4\times 4$ quadratic degenerate reaction-diffusion system.
@@
For the degenerate system (3), we will relate the entropy dissipation functional with the missing term $\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}$. In this article, for the degenerate case (3), we analyze the decay of entropy in dimension $N\leq 3$. We will establish our result only for dimension $N=3$. For dimension $N=1,2$, all the calculations are similar. In Proposition 9 We establish the following relation between the entropy dissipation and the missing term $\|\delta_{C}\|^{2}_{\mathrm{L}^{2}(\Omega)}$
$D(a,b,c)\geq\hat{C}(1+t)^{-\frac{1}{3}}\left(\|A-\overline{A}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|B-\overline{B}\|_{\mathrm{L}^{2}(\Omega)}^{2}+\|C-\overline{C}\|_{\mathrm{L}^{2}(\Omega)}^{2}\right).$
@@
In order to arrive at the above estimate we need a particular integral estimate of species $a$, which we will establish in Lemma 6. The particular relation is the following
$\|a\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\tilde{C}(1+t)^{\frac{1}{3}}\qquad\forall t\geq 0$
@@
where $\tilde{C}>0$ a constant, independent of time. This is the key estimate for the degenerate case (3). To establish this we will use a particular integral estimate in Lemma 5.
@@
These results, along with the Grönwall inequality, will lead us to the conclusion that the entropy functional decays sub-exponentially fast in time for both the degenerate systems (2) and (3). On the other hand, an application of the Cziszár-Kullback-Pinsker inequality yields the following lower bound on the relative entropy (see [4] for details):
$E(a,b,c)-E(a_{\infty},b_{\infty},c_{\infty})$ $\geq\frac{(3+2\sqrt{2})|\Omega|}{2M_{1}(9+2\sqrt{2})}\|a-a_{\infty}\|_{\mathrm{L}^{1}(\Omega)}^{2}$
$+$ $\frac{(3+2\sqrt{2})|\Omega|}{2M_{2}(9+2\sqrt{2})}\|b-b_{\infty}\|_{\mathrm{L}^{1}(\Omega)}^{2}+\frac{(3+2\sqrt{2})|\Omega|}{(M_{1}+M_{2})(9+2\sqrt{2})}\|c-c_{\infty}\|_{\mathrm{L}^{1}(\Omega)}^{2}.$
Combining all these estimates we arrive at our main results: Theorem 1 and Theorem 2.